OpenStudy (lovelyharmonics):

Eliminate the parameter. x = t2 + 1, y = t2 - 1

ganeshie8 (ganeshie8):

solve t2 from first equation, plug it in second equation

jimthompson5910 (jim_thompson5910):

Or you can subtract equation 2 from equation 1 x = t^2 + 1 y = t^2 - 1 --------- ??

OpenStudy (lovelyharmonics):

so t^2=-1 @ganeshie8 and that would = 0... @jim_thompson5910

jimthompson5910 (jim_thompson5910):

on the left side, you'll have x-y on the right you'll have (t^2 + 1)-(t^2 - 1) and the right side simplifies to t^2 - t^2 + 1 - (-1) = 0t^2 + 2 = 2

jimthompson5910 (jim_thompson5910):

So you'll get x-y = 2 using the subtraction method

OpenStudy (lovelyharmonics):

i only have these answers... y = x + 2, x ≥ 1 y = x - 2, x ≥ 1 y = x2 - 2, x ≥ 1 y = x2 + 2, x ≥ 1

jimthompson5910 (jim_thompson5910):

now solve x-y = 2 for y

OpenStudy (lovelyharmonics):

oh duh im stupid :P so y=x-2

jimthompson5910 (jim_thompson5910):

you're not stupid and yes, it's y=x-2

OpenStudy (lovelyharmonics):

thank you c:

jimthompson5910 (jim_thompson5910):

you're welcome